Circular No. 8692

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 F2 (CHRISTENSEN) E. J. Christensen reports the discovery of a comet on unfiltered CCD

images taken with the 1.5-m reflector in the course of the Mt. Lemmon Survey (discovery observation tabulated below). The Mar. 23 images show a diffuse, round 8′′ coma and no tail in four co-added 30-s images. Four co-added 60-s images in average seeing taken on Mar. 24.3 UT show a 10′′ moderately condensed coma, slightly elongated toward p.a. 320^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
Mar. 23.32673  12^{h}36^{m}46.^{s}72  −3^{o}57′46′′.6  19.7

The available astrometry (including prediscovery Spacewatch observations from Feb. 27), the following preliminary orbital elements, and an ephemeris appear on MPEC 2006-F29.

T = 2006 Mar. 17.736 TT  ω = 179.^{o}204
e = 0.65377  Ω = 8.258 }2000.0
q = 4.28917 AU  i = 20.413
a = 12.38814 AU n^{o} = 0.022604 P = 43.60 years

COMET 73P/SCHWASSMANN-WACHMANN The components ‘H’, ‘J’, and L (mentioned on IAUC 8685 on Mar.

8) were observed by P. Birtwhistle (Great Shefford, Berkshire, England) on Mar. 23.1 UT and by E. J. Christensen at Mt. Lemmon on Mar. 23.5. Birtwhistle gave the magnitudes as 18.9, 19.9, and 21.0, adding that component ‘H’ had a coma of diameter 9′′ and a tail 40′′ long in p.a. 255^{o}, that component ‘J’ had a coma of diameter 9′′ and a tail 20′′ long in p.a. 230^{o}, and that component ‘L’ had a very diffuse coma of diameter 5′′ and a possible tail 15′′ long in p.a. 245^{o}. Neither observer recorded component ‘K’, but both observers independently reported two new components, designated ‘M’ and ‘N’, estimated by Birtwhistle at mag 20.8 and 21.4, respectively, the former having a diameter of no more than 4′′ and no tail, and the latter having a stellar head and a faint tail 70′′ (possibly up to 90′′) long in p.a. 255^{o}. The astrometric observations are provided on MPEC 2006-F22. The new components ‘M’ and ‘N’ are in the vicinity of component ‘H’ and, using the earlier orbit (MPEC 2006-B27) for component ‘B’ with perihelion time adjusted as before, they correspond to T = 2006 June 8.28 and 8.30 TT, respectively, although component ‘N’ lies perhaps 40′′ north of the line of variation.

2006 March 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green